OpenStudy (anonymous):

4. Which value for x makes the sentence true? 4x - 3 = 21 x = 28 x = 6 x = 4 x = 24

pooja195 (pooja195):

add 3 to both sides

pooja195 (pooja195):

Remeber we are trying to ISOLATE X

OpenStudy (anonymous):

okay

OpenStudy (black_ninja123):

4*6=24 24-3=1 x=6 also the same thing as the last question

OpenStudy (anonymous):

7 AND 15

OpenStudy (lizz123):

4x - 3 = 21 4x + 3 = + 3 4x ________________ = ____________ 4x = 24 _________ _________ 4 4 x=6